# Fusing Dynamics Equation: A Social Opinions Prediction Algorithm with LLM-based Agents

### v1

@@ Abstract
DOE-LLM, LLM-based Agent, ABM, Social Opinions Prediction
@@ II Method
The DOE-LLM divides users into opinion leaders and opinion followers. Opinion leaders are constrained by a Cellular Automata (CA) model based on LLM role-playing, while opinion followers are governed by a dynamic system that combines the CA model and the SIR model.
@@ II-A1 Opinion Leaders: LLM with CA
$T_{ij}^{t}=\begin{cases}O_{j},&r=0\\ 0,&r=1\text{ or }\left|O_{j}^{t}-O_{i}^{t}\right|>\epsilon\\ \left(O_{j}^{t}-O_{i}^{t}\right)\cdot\sqrt{r\cdot S_{j}^{t}},&r\neq 1,0\text{ and }\left|O_{j}^{t}-O_{i}^{t}\right|\leq\epsilon\end{cases}$ (2)
@@ II-A2 Opinion Follower: CA with SIR
$O_{i}^{t+1}=O_{i}^{t+1}\cdot e^{-\gamma\cdot|O_{i}^{t+1}|}\quad\text{if}\quad\text{random}<\gamma$ (4)
@@
$O_{i}^{t+1}=\mathrm{clip}\left(\left(r\cdot O_{i}^{t}+w\cdot\frac{\sum_{j\in N_{i}}\left(O_{j}^{t}-O_{i}^{t}\right)\cdot\sqrt{r\cdot|O_{j}^{t}|}\cdot I\left(|O_{j}^{t}-O_{i}^{t}|\leq\epsilon\right)}{|N_{i}|}\right)\cdot e^{-\gamma\cdot|O_{i}^{t+1}|}\cdot I(\text{random}<\gamma),-1,1\right)$ (5)
Here $|N_{i}|$ is the number of neighbors around individual $i$, $|O_{j}^{t}-O_{i}^{t}|$ represents the opinion difference between neighbor $j$ and individual $i$, $I\left(|O_{j}^{t}-O_{i}^{t}|\leq\epsilon\right)$ is an indicator function that equals 1 if the opinion difference is within the threshold $\epsilon$, and 0 otherwise, and $\gamma$ is the decay rate indicating the natural decay of opinions.
@@ III-A Experimental Setup
[…]
@@ III-C Main Result
Our DOE-LLM model incorporates the SIR model to constrain the attitudes of opinion followers, successfully simulating the attitude loss characteristic of group attitudes. Compared to the LLM+ABM model, which focuses only on attitude changes due to news events, the DOE-LLM model enriches the natural attenuation of group attitudes on top of attitude changes, making the simulation of social news events more realistic. When using LLM alone, the attitude shows a sharp turn and remains stable at extreme values. When using the ABM model alone, good results are obtained in some events, but it lacks the interpretability of the DOE-LLM model. Detailed results are shown in Fig. 4 and Table. I.
The DOE-LLM model, through the action log of the LLM opinion leader part, can directly analyze group statements on a micro level, making it more valuable in predicting and controlling public opinion situations. We have discussed this issue in III-E.
@@ III-D Ablation Study
We discussed the performance of the DOE-LLM model when the CA constraint on core users is removed.
@@
DOE-LLM 0.3622 0.9653 0.3664 0.8950 0.3352 0.9605 0.3404 0.8842
DOE-LLM (Without CA) 0.5133 0.9489 0.7530 0.8671 0.7389 0.8952 0.5708 0.7367
Figure 5: DOE-LLM without CA
@@ IV Conclusion
In this paper, we optimized the model for news sentiment prediction by combining LLM and ABM, incorporating constraints from the CA and SIR models to preserve the natural attenuation of group attitudes under the guidance of LLM. Our proposed DOE-LLM divides users into opinion leaders and opinion followers. Opinion leaders are role-played based on LLM and constrained by the CA model, while opinion followers are part of a dynamic system that integrates the CA and SIR models. This innovative design significantly improves simulation accuracy and predictive efficiency.

### v3

@@ Abstract
FDE-LLM, LLM-based Agent, ABM, Social Opinions Prediction
@@ II Method
The FDE-LLM divides users into opinion leaders and opinion followers. Opinion leaders are constrained by a Cellular Automata (CA) model based on LLM role-playing, while opinion followers are governed by a dynamic system that combines the CA model and the SIR model.
@@ II-A1 Opinion Leaders: LLM with CA
$T_{ij}^{t}=\begin{cases}O_{j},&r=0\\ 0,&r=1\text{ or }\left|O_{j}^{t}-O_{i}^{t}\right|>\epsilon\\ \left(O_{j}^{t}-O_{i}^{t}\right)\cdot\sqrt{r\cdot|O_{j}^{t}|},&r\neq 1,0\text{and }\left|O_{j}^{t}-O_{i}^{t}\right|\leq\epsilon\end{cases}$ (2)
@@ II-A2 Opinion Follower: CA with SIR
$O_{i}^{t+1}=O_{i}^{t+1}\cdot e^{-\lambda\cdot|O_{i}^{t+1}|}\quad\text{if}\quad\text{random}<\gamma$ (4)
@@
$O_{i}^{t+1}=\mathrm{clip}\left(\left(r\cdot O_{i}^{t}+w\cdot\frac{\sum_{j\in N_{i}}\left(O_{j}^{t}-O_{i}^{t}\right)\cdot\sqrt{r\cdot|O_{j}^{t}|}\cdot I\left(|O_{j}^{t}-O_{i}^{t}|\leq\epsilon\right)}{|N_{i}|}\right)\cdot e^{-\lambda\cdot|O_{i}^{t+1}|}\cdot I(\text{random}<\gamma),-1,1\right)$ (5)
Here $|N_{i}|$ is the number of neighbors around individual $i$, $|O_{j}^{t}-O_{i}^{t}|$ represents the opinion difference between neighbor $j$ and individual $i$, $I\left(|O_{j}^{t}-O_{i}^{t}|\leq\epsilon\right)$ is an indicator function that equals 1 if the opinion difference is within the threshold $\epsilon$, and 0 otherwise, and $\lambda$ is the decay rate indicating the natural decay of opinions.
@@ III-A Experimental Setup
3. Parameter settings: Through experiments, it was found that we can use unified parameters for all models on similar reversal news. The parameters are obtained by grid search on the ABM algorithm to maximize the correlation coefficient. The specific settings are as follows: For the CA model: opinion resilience $r=0.99$, neighborhood influence coefficient $w=0.3$, opinion interaction threshold $\epsilon=0.5$; for the SIR model: infection rate $\beta=0.3$, recovery rate $\gamma=0.9$, decay rate $\lambda=0.5$.
@@ III-C Main Result
Our FDE-LLM model incorporates the SIR model to constrain the attitudes of opinion followers, successfully simulating the attitude loss characteristic of group attitudes. Compared to the LLM+ABM model, which focuses only on attitude changes due to news events, the FDE-LLM model enriches the natural attenuation of group attitudes on top of attitude changes, making the simulation of social news events more realistic. When using LLM alone, the attitude shows a sharp turn and remains stable at extreme values. When using the ABM model alone, good results are obtained in some events, but it lacks the interpretability of the FDE-LLM model. Detailed results are shown in Fig. 4 and Table. I.
The FDE-LLM model, through the action log of the LLM opinion leader part, can directly analyze group statements on a micro level, making it more valuable in predicting and controlling public opinion situations. We have discussed this issue in III-E.
@@ III-D Ablation Study
We discussed the performance of the FDE-LLM model when the CA constraint on core users is removed.
@@
FDE-LLM 0.3622 0.9653 0.3664 0.8950 0.3352 0.9605 0.3404 0.8842
FDE-LLM (Without CA) 0.5133 0.9489 0.7530 0.8671 0.7389 0.8952 0.5708 0.7367
Figure 5: FDE-LLM without CA
@@ IV Conclusion
In this paper, we optimized the model for news sentiment prediction by combining LLM and ABM, incorporating constraints from the CA and SIR models to preserve the natural attenuation of group attitudes under the guidance of LLM. Our proposed FDE-LLM divides users into opinion leaders and opinion followers. Opinion leaders are role-played based on LLM and constrained by the CA model, while opinion followers are part of a dynamic system that integrates the CA and SIR models. This innovative design significantly improves simulation accuracy and predictive efficiency.